OpenStudy (anonymous):

2. –4x + 3y = –12 –2x + 3y = –18 (1 point)

OpenStudy (anonymous):

What exactly is the question?

OpenStudy (anonymous):

what is the solution of the system? Use a graph.

OpenStudy (anonymous):

i have the graphs its a multiple choice @cassidyohrt

OpenStudy (anonymous):

-4x + 3y = -12 --->(-1) -2x + 3y = -18 ---------------- 4x - 3y = 12 (result of multiplying by -1) -2x + 3y = -18 ---------------add 2x = - 6 x = -6/2 = -3 now sub in -3 for x -2x + 3y = -18 -2(-3) + 3y = -18 6 + 3y = -18 3y = -18 - 6 3y = - 24 y = - 8 check... -4x + 3y = -12 -4(-3) + 3(-8) = -12 12 - 24 = -12 -12 = -12 (correct) x = -3 and y = -8

OpenStudy (anonymous):

thanks @kelliegirl33

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

and @sweetburger